OpenStudy (anonymous):

let G be group such that (a*b)^2 = a^2 * b^2 for a,b element G, show that is abelian.

OpenStudy (anonymous):

You know that \[(G,*)\text{ is a group. To show that it's abelian, you must show commutativity,}\\ \text{i.e. }\;\;a*b=b*a\]

OpenStudy (anonymous):

d you mean that i must say (a * b)^2 = a^2 * b^2 = b^2 * a^2 = ( b * a)^2 thus <G , * is abeian

OpenStudy (anonymous):

Is the operation you're using here multiplication? Or just the general "star"? If it's the second, I don't think you can make that conclusion right away.

OpenStudy (anonymous):

the oparation is just general star

OpenStudy (anonymous):

You seem to have treated * as multiplication, which doesn't work. For this particular operation, I'm not sure how to show whether it's commutative in this case. It's the (a*b)^2 that's bothering me.